OpenStudy (anonymous):

Help Help Help?!?!? What is the equation of the line of symmetry for the graph of y = 2(x + 3)2 – 4? Answer x = –4 x = –2 x = 3 x = –3

OpenStudy (anonymous):

@ganeshie8 can you help?

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

are you sure its not -3

ganeshie8 (ganeshie8):

it is x = -3, as the line of symmetry passes through vertex for a parabola

OpenStudy (anonymous):

\[2(x+3)^{2}-4=2x ^{2}+12x+18-4\]intersecting with y=14,\[2x ^{2}+12x=0\rightarrow x ^{2}+6x=0\rightarrow x=0,x+6=0\rightarrow x=0,x=-6\]Thus your simetry will be at x=-3 PS: my bad you were right, got hasty i guess...